# Statistical Taylor Expansion: A New and Path-Independent Method for Uncertainty Analysis 111

### v18

@@ 7 Mathematical Library Functions
The reasons why $(x^{p})^{1/p}-x$ has larger value errors than $\log(e^{x})-x$, and $e^{\log(x)}-x$ has nearly no value error, are not clear. It could relate to floating-point representation which stores each value $x$ as $\log_{2}(x)$. Figure 5 also suggests that using linear approximation $e^{x\pm\delta x}$ and $\log(x\pm\delta x)$ have less error than $(x\pm\delta x)^{p}$.
@@ 9.4 Trigonometric Library Errors
  • Figure 23 shows that the absolute value difference between the Library $\sin(x)$ and the Quart $\sin(x)$ increases approximately linearly with $|x|$.
@@ 9.5 Using Quart Sine for Sin/Cos Signals
When $L=18$ and $\delta x=0$, Figure 25 shows that the error distributions of Sin/Cos signal resemble Normal distributions, with an additional Delta-like distribution at $\tilde{z}=0$. The error distribution of the reverse transformation is structured on top of the Normal distribution, suggesting that the reverse transformation is more sensitive to numerical errors in the sine function.
@@ 9.6 Using Library Sine for Sin/Cos Signals
When $\delta x=0$, using the Library sine function, for a sine wave with a frequency of $3$, Figures 27, 28, and 29 present the value errors for the forward, reverse, and round-trip transformations, respectively. In the reverse transformation, the value errors exhibit a clear trend of increasing with the time index. These large value errors appear systematic rather than random and visually resemble a resonant pattern. Similar increases are observed at other frequencies and FFT orders, as well as in computational results obtained using mathematical libraries such as SciPy. In contrast, such resonance is absent from the round-trip transformation shown in Figure 29, as well as when using the Quart sine function. Figure 30 demonstrates that the error deviations increase with sine or cosine frequency, regardless of FFT order $L$ when $L>8$. Figure 23 indicates that the numerical errors of $cos(x)/sin(x)$ obtained using the Library sine function increase with a periodicity of $\pi$, which can resonate with a signal whose periodicity is an integer multiply of $\pi$, producing the resonant pattern shown in Figure 28. At higher frequency, the resonant beats between the signal and the numerical errors in the Library sine function become stronger. To suppress this numerical error resonance, an input noise of $\delta x=10^{-14}$ must be added to the sine or cosine signals. Such resonance of numerical errors can easily and mistakenly be taken as signals.
@@ 9.9 Ideal Coverage
Adding noise to the input can suppress unspecified input errors. After adding a Gaussian input noise of $\delta x=10^{-3}$ to a Linear signal when using the Library sine function, the error distributions for both the forward and the reverse transformations become Normal, whereas the error distribution for the round-trip transformation becomes Delta. Figure 38 illustrates the corresponding error deviations and uncertainty deviations versus FFT order:
@@ 11.2 Improvements Needed
Bound momentum $\zeta(n,\kappa)$ should be extended to all probability distributions. The choice of ideal bounding range $\hat{\kappa}$ should be extended to other distributions, perhaps by using the same value of $1-\zeta(2,\kappa)$. The procedure to find the bound moments $\kappa$ for each sample count $N$ should be developed for discrete distributions. The statistical meaning for ideal leakage $\varepsilon\in[0,1]$ in the context of its choice of ideal bounding range $\hat{\kappa}$ needs further investigation: Is $\varepsilon$ the possibility for the result to be not bound by $\hat{\kappa}$?

### v19

@@ 7 Mathematical Library Functions
The reasons why $(x^{p})^{1/p}-x$ has larger value errors than $\log(e^{x})-x$, and $e^{\log(x)}-x$ has nearly no value error, are not clear. It could relate to floating-point representation. Figure 5 also suggests that using linear approximation $e^{x\pm\delta x}$ and $\log(x\pm\delta x)$ have less error than $(x\pm\delta x)^{p}$.
@@ 9.4 Trigonometric Library Errors
  • Figure 23 shows that the value difference between the Library $\sin(x)$ and the Quart $\sin(x)$ increases approximately linearly with $|x|$.
@@ 9.5 Using Quart Sine for Sin/Cos Signals
When $L=18$ and $\delta x=0$, Figure 25 shows that the error distributions of Sin/Cos signal resemble Normal distributions, with an additional Delta-like distribution at $\tilde{z}=0$ for the forward transformation. The error distribution of the reverse transformation is structured on top of the Normal distribution, suggesting that the reverse transformation is more sensitive to numerical errors in the sine function.
@@ 9.6 Using Library Sine for Sin/Cos Signals
When $\delta x=0$, using the Library sine function, for a sine wave with a frequency of $3$, Figures 27, 28, and 29 present the value errors for the forward, reverse, and round-trip transformations, respectively. In the reverse transformation, the value errors exhibit a clear trend of increasing with the time index. These large value errors appear systematic rather than random and visually resemble a resonant pattern. Similar increases are observed at other frequencies and FFT orders, as well as in computational results obtained using mathematical libraries such as SciPy. In contrast, such resonance is absent from the round-trip transformation shown in Figure 29, as well as when using the Quart sine function. Figure 30 demonstrates that the error deviations increase with sine or cosine frequency, regardless of FFT order $L$ when $L>8$. Figure 23 indicates that the numerical errors of $sin(x)$ obtained using the Library sine function increase with a periodicity of $\pi$, which can resonate with a signal whose periodicity is an integer multiply of $\pi$, producing the resonant pattern shown in Figure 28. At higher frequency, the resonant beats between the signal and the numerical errors in the Library sine function become stronger. To suppress this numerical error resonance, an input noise of $\delta x=10^{-14}$ must be added to the sine or cosine signals. Such resonance of numerical errors can easily and mistakenly be taken as signals.
@@ 9.9 Ideal Coverage
Figure 38: Error distributions for Linear signals with $10^{-3}$ input noise for forward, reverse and roundtrip transformations (legend) computed using the Library sine function. The input noise is either Gaussian or white (legend). The FFT order is $18$.
@@
Adding noise to the input can suppress unspecified input errors. After adding a Gaussian input noise of $\delta x=10^{-3}$ to a Linear signal when using the Library sine function, the error distributions for both the forward and the reverse transformations become Normal, whereas the error distribution for the round-trip transformation becomes Delta, as shown in Figure 38. Figure 39 illustrates the corresponding error deviations and uncertainty deviations versus FFT order:
@@ 11.2 Improvements Needed
Bound momentum $\zeta(n,\kappa)$ should be extended to all probability distributions. The choice of ideal bounding range $\hat{\kappa}$ should be extended to other distributions, perhaps by using the same value of linear leakage $\varepsilon\equiv 1-\zeta(2,\kappa)$. The procedure to find the bound moments $\kappa$ for each sample count $N$ should be developed for discrete distributions. The statistical meaning for ideal leakage $\varepsilon\in[0,1]$ in the context of its choice of ideal bounding range $\hat{\kappa}$ needs further investigation: Is $\varepsilon$ the possibility for the result to be not bound by $\hat{\kappa}$?